A USA multinational is considering a European opportunity. The multinational is considering an initial investment of €12,000. The size and timing of the before interest and taxes cash flows is as follows:

Year

1

2

3

4

5

EBIT (€)

800

600

700

1000

650

The following information is provided:

• Inflation in US, πₛ = 4.2%

• Inflation in the Eurozone, π€ = 7.2%

• Corporate tax rate in Eurozone, T = 40%

• Cost of Equity to an unlevered USA firm, Kud = i$ = 15.2%

• Current spot exchange between the US dollar and the Euro is Sₒ($/€) = $1.31/€

Required:

Using NPV method, appraise this project from the USA shareholders perspective and advise them accordingly
Answer #1
Year Cash Flow in Euro Post Tax Income (1-tax rate) in Euro Dollar/Euro Cashflow in USD Present Value
0 -12000 -12000 1.310 -9160.31 -9160.31
1 800 480 1.348 356.16 309.17
2 600 360 1.387 259.64 195.65
3 700 420 1.426 294.44 192.59
4 1000 600 1.468 408.86 232.15
5 650 390 1.510 258.32 127.32
NPV -8103.44

Here cash flow in Euro is post-tax income in Euro is = Cash Flow * ( 1- tax rate)

Year 01    = 800*(1-40%) = 480

Year 02 = 600*(1-40%) = 360  And so on...

Expected Dollar Value in Next Year = Current Dollar/Euro Rate * ( 1+Eurozone Inflation)/(1 + US $ Inflation)

   Year 01 Dollar Value = 1.310 * (1.072/1.042) = 1.348

Year 02 Dollar Value = Year 01 Dollar Value * (1.072/1.042) = 1.348 * (1.072/1.042) = 1.387 And So on...

Cash Flow in USD = Post Tax Income in Euro / Dollar Euro Conversion  

Year 00  Cash Flow = -12000/1.310 = -9160.31

Year 01 Cash Flow = 480/1.348 = 356.16 And So on....

Present Value of Dollar

Year 01 $ Value = Net cashflow in USD / (1+Cost of Capital)

= 356.16 / (1 + 15.2%) = 309.17

Year 02 $ Value =  Net cashflow in USD /(1+Cost of Capital)^2

= 259.64 / (1 + 15.2%)^2  = 195.65 and So on....

Here Net Present Value of the Project is -Negative. With an investment of around 9160 USD Return will be around 1057 USD in present value.

So, Share holders should Not invest in this project.